# Cross-Category Functional Grasp Tansfer

### v2

@@ III-D Grasp Synthesis based on Transferred Touch Code
pan trigger_sprayer pliers stapler Avg flashlight scrather electric_teapot twohead_wrench spatula paint_roller Avg
1 No 90.88% 43.73% 23.57% 45.63% 58.59% 54.98% 59.88% 43.35% 54.09% 68.15% 42.65% 49.53%
2 Shape 97.11% 57.04% 45.12% 65.10% 68.87% 61.86% 65.76% 53.52% 66.21% 78.26% 59.21% 60.97%
3 Layout 94.07% 51.59% 45.43% 60.74% 65.23% 61.05% 64.15% 47.80% 71.82% 75.02% 57.07% 57.85%
4 Grasp type 95.87% 56.31% 39.53% 62.80% 67.53% 61.49% 65.09% 49.39% 55.28% 76.60% 62.36% 59.15%
5 S+L 97.36% 54.32% 45.62% 64.91% 70.32% 64.15% 66.10% 51.65% 84.26% 81.65% 62.54% 63.16%
6 S+G 97.22% 56.14% 45.61% 66.84% 71.93% 63.38% 68.01% 59.43% 78.85% 81.76% 71.61% 63.09%
7 L+G 97.41% 55.90% 49.48% 69.13% 73.20% 67.15% 68.92% 57.92% 73.78% 80.37% 73.31% 64.11%
8 S+L+G 97.56% 57.77% 45.15% 67.30% 73.89% 66.56% 71.93% 60.66% 88.53% 81.38% 77.15% 64.88%
9 PointNet 96.63% 55.68% 42.13% 63.21% 68.15% 63.65% 67.30% 56.03% 85.72% 77.93% 72.03% 60.43%
10 PointNet++ 97.33% 55.39% 43.53% 65.19% 70.89% 65.11% 68.96% 58.26% 83.66% 79.25% 75.87% 62.01%
@@
Our grasp optimization is mainly divided into two stages. The first stage only adjusts the translation and rotation of the hand, with an initial learning rate of 0.001 and a total of 100 rounds of optimization. The second stage mainly adjusts the joint angles of the dexterous hand, with an initial learning rate of 0.001, while the learning rates for translation and rotation are reduced to 0.0001, with a total of 200 optimization rounds. Thanks to the functional grasp types we proposed, this letter synthesizes higher-quality functional grasps using only classical optimization methods compared to existing network-based approaches [5]. The results of grasp synthesis are presented in Section IV-B.
@@ IV-A Result of Similarity-based Touch Code Transfer
As described in Sec. III-C, our framework relies on three types of inter-object similarity to transfer the touch code of the training objects to new objects. To evaluate the effect of the touch code prediction, we tested a total of 49 object categories, including 12 unseen categories used to evaluate the cross-category prediction performance. Fig. 8 shows some of the test results, with seen categories on the left of the dashed line and the unseen categories on the right, where each object displays the touch code label (left) and prediction result (right) respectively. It can be observed that our framework not only predicts the touch codes of known categories effectively but also copes well with new categories. For example, the flashlight class highlighted in the red box was not involved in either the auto-encoder training or the GAT training, but associated with objects like electric shavers and electric toothbrushes through layout and grasp type similarities, enabling the flashlight to successfully achieve the touch code prediction of the head, switch, and body, without being misjudged as a hammer.
@@
Moreover, Tab. I presents the quantitative results for touch code prediction, using the evaluation metric of mean intersection over union (mIoU). ‘Row-1’ represents the results of the baseline without using inter-object relationships, while ‘Row-8’ represents the results of our transfer framework when both training and testing use correct shape, layout, and grasp type similarities. By comparing ‘Row-1’ and ‘Row-8’, it is evident that in both seen and unseen object categories, our prediction framework shows significant improvement than the baseline. We believe that when similarities are not used to associate objects, the network can only rely on its memory to handle some simple objects, such as 90.88% mIoU of the pans in ‘Row-1’. However, it is incapable of handling more complex objects, such as trigger sprayers, pliers, and flashlights, whose mIoUs of ‘Row-1’(43.73%, 23.57%, 54.98%) are significantly lower than ‘Row-8’(57.77%, 45.15%, 66.56%). In addition, based on the well-trained model of ‘Row-8’, we also evaluated the prediction effect when building test graphs based on PointNet and PointNet++ respectively, as shown in ‘Row-9’ and ‘Row-10’ (where the prediction accuracies for shape, layout, and grasp type similarity of PointNet and PointNet++ are 90.70%, 89.74%, 91.55% and 91.36%, 91.07%, 92.97%, respectively). As shown in Tab. I, the results of ‘Row-9’ and ‘Row-10’ are slightly inferior to ‘Row-8’.
@@
Ablation Studies I . We conducted ablation studies on three similarities (rows 2-7), and some interesting conclusions can be drawn from these results: 1) Based on Avg mIoU of both seen and unseen categories, the results can be roughly divided into three groups, where the multi-relations groups (rows 5,6,7,8) outperform the single-relation groups (rows 2,3,4), while the no-relation group (row 1) has a significant gap with the other two groups. Therefore, it can be concluded that good knowledge transfer requires sufficient associations between objects. 2) The performance of ‘L+G’ (row 7) is better than ‘S+L+G’ (row 8) in some objects, such as the flashlight. We speculate that for flashlights, the association based on layout and grasp type similarity with objects like electric shavers is more crucial. In this case, introducing associations based on shape similarity with objects like hammers might lead to interference. Nevertheless, this does not imply that shape similarity is not important, as adding shape similarity for most objects can improve accuracy. This inspires us to explore how to intelligently adjust associations between objects to improve performance next.
@@
Ablation Studies II. Although the above experiments demonstrate the effectiveness of our framework in predicting touch code, it may be due to the network has remembered the mapping rules from object initial features to the touch code features, rather than relying on associations to transfer knowledge from previously learned samples to new objects. Therefore, we conducted the following two ablation experiments. 1) In our inference stage, the test object is connected to the training graph. To verify whether the learned objects in the training graph are necessary for predicting touch code, we generated a 360-node test graph containing only test objects, without using training graph as the connection. The prediction result is shown in Fig. 9(a), which reveals that the network is completely unable to handle these test samples, and even fails to retain the object structural information contained in the initial feature. 2) The second is to verify whether the similarities we proposed are effective in predicting the touch code. To achieve this, we randomly connect the test object to the previously learned objects in the training graph instead of using the correct associations. Fig. 9(b) shows the results. Taking the first mouse as an example, its result is completely incorrect. The reason is that the mouse is mistakenly associated with objects like guns and drills, leading to incorrect touch code prediction. The above two experiments illustrate that our prediction framework does not simply memorize the mapping between features, but actually achieves human-like reasoning, in which both reasonable inter-object associations and previously learned samples are crucial.
@@ IV-B Functional Grasp Synthesis
As mentioned in Section III-D, we employ an optimization method based on predicted touch code and functional grasp type to synthesis functional grasp. It can be observed from the results in Fig. 10 that objects from both seen and unseen categories can achieve functional grasps effectively. For example, within the red box, the dexterous hand’s index and middle fingertip accurately land on the left and right buttons of the mouse, while the thumb, ring finger, and little finger grasp the sides of the mouse. Another example is the electric teapot in the blue box, where the thumb presses the switch above the handle, and the remaining fingers hold the handle. Moreover, we use Isaac Gym [35] to calculate the success rate of grasps, where the object is considered to be grasped successfully if it does not fall within 10 seconds under random external force interference after being grasped. As indicated in the Tab. II, our method demonstrated a good success rate on both seen and unseen objects. For instance, in the case of cross-category grasp synthesis, the success rate for flashlights reached 83.3%.
@@
drill mouse gun sprayer axe electric_teapot spoon flashlight
Ours 85.7 70.0 63.2 76.5 66.7 66.7 75.0 83.3
Figure 11: Some grasp results of existing works. Left: [6]. Right: [36].
In addition, it is not difficult to observe from Fig. 10 that our grasp poses are highly reasonable, approaching human-like grasps. This is not only attributed to the precise touch code prediction mentioned earlier but also benefited from the functional grasp types we proposed. Our functional grasp types can not only used to build cross-category associations for touch code transfer, but also provide the reasonable initialization for dexterous hand in functional grasp synthesis. This enables the dexterous hand to make correct contact with the object in the right pose, as illustrated in the grasp of the second gun in Fig. 10, where the index finger accurately presses the trigger inside the loop. This advantage can be demonstrated through a comparison with two existing touch code-based functional grasp generation works  [6, 36] that have not used our functional grasp types. As shown in Fig. 11, the grasping results of these two works exhibit unreasonable grasping poses, such as the grasp of the first gun, where the dexterous hand adopt a strange pose to strive for the correct contact, and the index finger approaches the trigger in the opposite direction.
@@ IV-C Comparison with Existing Works
In order to achieve cross-category grasp transfer for multi-fingered dexterous hands, some methods [1, 11] propose leveraging the local similarity of objects. However, these methods lack consideration for the overall structure and layout of objects, and can only achieve transfer of stable grasps for dexterous hands, rather than functional grasps needed for precise manipulation of objects such as flashlights and guns. The most relevant work to ours is [5], where they achieve functional grasp transfer within a category based on dense correspondences of surface points, achieving an 83.8% grasp success rate in the flashlight category. In comparison, this letter breaks through the intra-category constraints and achieves cross-category functional grasp transfer, with a success rate of 83.3% in the flashlight category.
@@ V CONCLUSIONS
To transfer the available functional grasps of objects to the other categories of objects, this letter analyzes the behavior of human grasping across categories and proposes three similarity relationships related to the grasp transfer. We propose the knowledge graph-based touch code transfer and optimization-based grasping synthesis framework, achieving cross-category functional grasp synthesis. It can be said that this letter creatively provides new insights and new benchmark for cross-category functional grasp transfer. However, for the convenience of reader comprehension and method implementation, this letter employs relatively basic models. Therefore, there is still considerable room for improvement in performance. In the future, we will improve the network to enhance its performance and design a more flexible structure to further achieve grasping across tasks. By dynamically adjusting the links between nodes, we aim to enable objects to have multiple uses for different tasks.

### v4

@@ III-D Grasp Synthesis based on Transferred Touch Code
hammer pan gun mug Avg flashlight scrather electric_teapot trigger_sprayer spatula paint_roller Avg
1 No 66.66% 90.16% 40.19% 82.52% 59.28% 51.03% 60.28% 40.37% 12.25% 68.55% 40.12% 44.34%
2 Shape 71.55% 96.88% 51.81% 88.42% 68.96% 61.33% 65.85% 52.30% 35.15% 78.64% 58.07% 58.29%
3 Layout 70.26% 94.37% 47.07% 87.96% 66.11% 60.18% 64.25% 46.02% 39.34% 75.20% 56.89% 55.28%
4 Grasp type 69.89% 95.09% 51.76% 86.01% 67.92% 61.46% 65.39% 48.21% 37.83% 76.66% 62.08% 57.92%
5 S+L 72.29% 97.43% 56.74% 92.72% 70.60% 63.83% 66.42% 50.82% 42.18% 81.19% 62.14% 62.23%
6 S+G 72.57% 97.01% 53.76% 92.28% 72.69% 63.50% 68.58% 59.17% 40.70% 81.84% 71.05% 62.47%
7 L+G 73.58% 97.25% 57.56% 92.16% 72.83% 67.02% 69.18% 58.22% 41.91% 80.46% 72.87% 63.55%
8 S+L+G 75.75% 97.52% 57.47% 92.92% 73.49% 66.96% 72.33% 60.36% 45.36% 81.19% 76.83% 64.02%
9 PointNet 70.71% 95.51% 41.26% 92.13% 68.93% 62.13% 67.67% 55.62% 42.39% 77.96% 72.33% 59.32%
10 PointNet++ 71.24% 96.15% 42.37% 92.25% 71.57% 63.56% 69.34% 57.73% 43.11% 79.17% 73.98% 61.66%
@@
To synthesize the functional grasp of an object, we first provide the optimization algorithm with a good initial pose of Shadowhand based on our functional grasp type ($G$). Specifically, we annotate nine pre-grasp poses for Shadowhand corresponding to nine functional grasp types. We then train the PointNet++ to predict the grasp type of the target object, thereby obtaining its pre-grasp pose. To make the touch code ($O_{c}$) applicable to Shadowhand, we select 16 links of Shadowhand corresponding to 16 parts of human hand as shown in Fig. 6. After completing the above steps, we refer to the loss functions from prior work [5] and use three objective functions to optimize the functional grasp of the Shadowhand. The first is the attraction function, which minimizes the distance between the links of the dexterous hand and the related object area specified by the touch code (the area with touch code 1), such as the distance between the distal thumb and the flashlight switch. The second is the repulsion function, which serves to keep the links of the dexterous hand away from non-related object areas specified by the touch code (the area with touch code 0). The third is the hand regularization function, which consists of two parts. The first part keeps the joint angles of the dexterous hand within a reasonable range, and the second part prevents the fingers of the dexterous hand from colliding with each other. For detailed formulas, please refer to our prior work [5].
Our grasp optimization is divided into two stages. The first stage only adjusts the translation and rotation of the hand, with an initial learning rate of 0.001 and a total of 100 rounds of optimization. The second stage mainly adjusts the joint angles of the dexterous hand, with an initial learning rate of 0.001, while the learning rates for translation and rotation are reduced to 0.0001, with a total of 200 optimization rounds. Thanks to the functional grasp types we proposed, this letter synthesizes higher-quality functional grasps using only classical optimization methods compared to existing network-based approaches [5].
@@ IV-A Result of Similarity-based Touch Code Transfer
As described in Sec. III-C, our framework relies on three types of inter-object similarity to transfer the touch code of the training objects to new objects. We test a total of 49 object categories, including 15 unseen categories used to evaluate the cross-category touch code prediction performance. Fig. 7 shows some of the test results, with seen categories on the left of the dashed line and the unseen categories on the right, where each object displays the touch code label (left) and prediction result (right) respectively. It can be observed that our framework not only predicts the touch codes of known categories effectively but also copes well with new categories. For example, the flashlight class highlighted in the red box is not involved in either the auto-encoder training or the GAT training, but associated with objects like electric shavers and electric toothbrushes through layout and grasp type similarities, enabling the flashlight to successfully achieve the touch code prediction of the head, switch, and body, without being misjudged as a hammer.
@@
Moreover, Tab. I presents the quantitative results for touch code prediction, using the evaluation metric of mean intersection over union (mIoU). ‘Row-1’ represents the results of the baseline without using inter-object relationships, ‘Row-8’ represents the results of our transfer framework when both training and testing use the correct shape, layout, and grasp type similarities. By comparing ‘Row-1’ and ‘Row-8’, it is evident that in both seen and unseen object categories, our method shows significant improvement over the baseline. We believe that when similarities are not used to associate objects, the network can only rely on its memory to handle some simple objects, such as 90.16% mIoU of the pans in ‘Row-1’. However, it is incapable of handling more complex objects, such as guns, flashlights and trigger sprayers, whose mIoUs of ‘Row-1’(40.19%, 51.03%, 12.25%) are significantly lower than ‘Row-8’(57.47%, 66.96%, 45.36%). In addition, based on the well-trained model of ‘Row-8’, we also evaluate the prediction effect when constructing test graphs using PointNet or PointNet++, where the shape, layout and grasp type similarities between the test object and known objects are predicted by PointNet or PointNet++. As shown in Tab. I, the results of ‘Row-9’ and ‘Row-10’ are slightly inferior to ‘Row-8’.
@@
Ablation Studies I . We conduct ablation studies on three similarities (rows 2-8), and some interesting conclusions can be drawn from these results: 1) Based on Avg mIoU of both seen and unseen categories, the results can be roughly divided into three groups, where the multi-relations groups (rows 5-8) outperform the single-relation groups (rows 2-4), while the no-relation group (row 1) has a significant gap with the other two groups. Therefore, it can be concluded that good knowledge transfer requires sufficient associations between objects. 2) The performance of ‘L+G’ (row 7) is better than ‘S+L+G’ (row 8) in some objects, such as the flashlight. We speculate that for flashlights, the association based on layout and grasp type similarity with objects like electric shavers is more crucial. In this case, introducing associations based on shape similarity with objects like hammers might lead to interference. Nevertheless, it does not imply that shape similarity is not important, as adding shape similarity for most objects can improve accuracy. This inspires us to explore how to intelligently adjust associations between objects to improve performance next.
@@
Ablation Studies II. In our inference stage, the test object is connected to the training graph. To verify whether the training graph (i.e., learned objects) are necessary for predicting touch code, we generate a 360-node test graph containing only test objects. Prediction results are shown in Fig. 9(a), which reveal that the network is completely unable to handle these test samples, and even fail to retain the object structural information. Additionally, to further validate whether the similarities we proposed are effective in transferring the touch code, we randomly connect the test object to the previously learned objects in the training graph instead of using the correct associations, and the results are shown in Fig. 9(b). Taking the first mouse as an example, its result is completely incorrect due to a mistaken association with objects like guns and drills, leading to incorrect touch code transfer. The above two experiments illustrate that our framework does not simply memorize the mapping between features, but actually achieves human-like reasoning, in which both reasonable inter-object associations and previously learned samples are crucial.
@@ IV-B Functional Grasp Synthesis
As mentioned in Sec. III-D, we employ an optimization method based on predicted touch code and functional grasp type to synthesise functional grasp. It can be observed from the results in Fig. 8 that objects from both seen and unseen categories can achieve functional grasps effectively. For example, within the red box, the dexterous hand’s index and middle fingertip accurately land on the left and right buttons of the mouse, while the thumb, ring finger, and little finger grasp the sides of the mouse. Another example is the electric teapot in the blue box, where the thumb presses the switch above the handle, and the remaining fingers hold the handle. These successful functional grasp results stem from accurate touch code predictions and the effective pre-grasp poses provided by our functional grasp types. Moreover, we use Isaac Gym [30] to calculate grasp success rate. A successful grasp must lift the object against gravity for 5s and then withstand an additional 5s of continuously and randomly changing force disturbances. As indicated in the Tab. II, our method demonstrate a good success rate on both seen and unseen objects. For instance, in the case of cross-category grasp synthesis, the success rate for flashlights reached 83.3%.
@@
bottle knife mouse gun drill electric_teapot sprayer flashlight
Ours 67.9 81.8 70.0 63.2 85.7 66.7 70.6 83.3
TABLE III: Comparison with Existing Works.
Method Grasp Success Rate
Drill Spray Bottle Flashlight Bottle Knife
 [13] 58.3% 57.1% - - -
 [5] - - 83.8% 71.7% 90.5%
Ours 85.7% 70.6% 83.3% 67.9% 81.8%
@@ IV-C Comparison with Existing Works
To the best of our knowledge, the most relevant works to ours are [5] and [13], both of which aim to achieve functional grasp transfer within a single category. Specifically, [13] transfers grasp poses from one sample to other samples within the same category based on shape similarity, achieving grasp success rates of 58.3% and 57.1% on drills and spray bottles, respectively. [5] transfers touch codes to other similar samples by densely matching surface points of objects within the same category, achieving success rates of 83.8% on flashlights, 71.7% on bottles, and 90.5% on knives. Both methods require separate network to be trained for each object category. In contrast, our method is not limited to within-category transfer but instead uses a unified network model to achieve functional grasp transfer across different categories. In terms of grasp success rates, as shown in Tab. III, our cross-category transfer method performs on par with in-class transfer methods, with some categories even surpassing them. For instance, the results for the drill and sprayer categories exceed those results in [13]. In terms of inference time, [13] requires an average of 7s, and [5] takes 1.5s. In contrast, our method only needs 0.81s, with 0.05s for data processing, 0.16s for touch code transfer, and an average of 0.65s for grasp generation, demonstrating superior real-time performance.
@@ V CONCLUSIONS
To transfer the available functional grasps of objects to the other categories of objects, this letter analyzes the behavior of human grasping across categories and proposes three similarity relationships related to the grasp transfer. We propose the knowledge graph-based touch code transfer and optimization-based grasping synthesis framework, achieving cross-category functional grasp synthesis. It can be said that this letter creatively provides a new insights and new benchmark for cross-category functional grasp transfer. However, for the convenience of reader comprehension and method implementation, this letter employs relatively basic models. Therefore, there is still considerable room for improvement in performance. In the future, we will improve the network to enhance its performance and design a more flexible structure to further achieve grasping across tasks. By dynamically adjusting the links between nodes, we aim to enable objects to have multiple uses for different tasks.